# Scaling Towards the Information Boundary of Instruction Set: InfinityInstruct-Subject Technical Report

### v1

@@ 1 Abstract
Instruction tuning has become a foundation for unlocking the capabilities of large-scale pretrained models and improving their performance on complex tasks. Thus, the construction of high-quality instruction datasets is crucial for enhancing model performance and generalizability. Although current instruction datasets have reached tens of millions of samples, models finetuned on them may still struggle with complex instruction following and tasks in rare domains. This is primarily due to limited expansion in both “coverage” (coverage of task types and knowledge areas) and “depth” (instruction complexity) of the instruction set. To address this issue, we propose a systematic instruction data construction framework, which integrates a hierarchical labeling system, an informative seed selection algorithm, an evolutionary data synthesis process, and a model deficiency diagnosis with targeted data generation. These components form an iterative closed-loop to continuously enhance the coverage and depth of instruction data. Based on this framework, we construct InfinityInstruct-Subject, a high-quality dataset containing $\sim$1.5 million instructions. Experiments on multiple foundation models and benchmark tasks demonstrate its effectiveness in improving instruction-following capabilities. Further analyses suggest that InfinityInstruct-Subject shows enlarged coverage and depth compared to comparable synthesized instruction datasets. Our work lays a theoretical and practical foundation for the efficient, continuous evolution of instruction datasets, moving from data quantity expansion to qualitative improvement.
@@ 2 Introduction
Instruction tuning serves as a cornerstone for unleashing the vast knowledge and reasoning capabilities of large pretrained models (Longpre et al., 2023; Ouyang et al., 2022). Consequently, the construction of high-quality instruction datasets has become essential for improving model performance and generalization ability. Several efforts have constructed instruction datasets through manual annotation or automatic synthesis, with the total size of these datasets having reached tens of millions (Ahmad et al., 2025; Nayak et al., 2024). However, current large models still struggle with complex instruction-following tasks and show limitation in “rare” tasks with low frequency (Qin et al., 2025).
A main reason would be that existing works overlooked the distribution of existing data, and thus failed to efficiently enhance the “coverage and “depth of synthesized instructions, and thus limit the model performance in solving rare tasks and difficult tasks (Wang et al., 2024; Liu et al., 2024; Long et al., 2024). The “coverage of an instruction dataset refers to the coverage of task types and knowledge categories it covers, while the ‘depth reflects the complexity of instructions within individual tasks. Both two dimensions are essential to enlarge the capability boundaries of large models (Shengyu et al., 2023). Expanding the coverage improves the model’s ability to generalize across domains, while increasing depth helps enhance complex reasoning ability and compositional generalizability (Sinha et al., 2024; Zhong et al., 2025). However, a key challenge remains: how to design effective, controllable and interpretable instruction generation strategies that continuously expand the instruction space along both dimensions, guiding models to acquire higher-level capabilities (Zhang et al., 2023; Qian et al., 2022; Qin et al., 2022). Currently, there is a lack of unified methodology and empirical research in this area.
@@
Based on this framework, we construct the InfinityInstruct-Subject (InfInstruct-Sub) dataset, which contains $\sim$1.5 million high-quality instructions. The code and dataset are available at https://github.com/BAAI-DIPL/InfinityInstruct-Sub and https://huggingface.co/datasets/BAAI/Infinity-Instruct/tree/main/Gen. Evaluations on multiple benchmarks demonstrate its effectiveness in improving instruction-following ability. Across foundation models, our tuned models outperform their official instruction-tuned counterparts. Data analysis further reveals that, compared with comparable synthetic datasets such as Magpie (Wei et al., 2023) and UltraChat(Ding et al., 2023), our framework could obtain an instruction set with higher coverage and complexity. Moreover, interestingly, we observe a scaling law in the co-occurrence structure of tags of instructions: the co-occurrence frequency of a tag with others follows a negative logarithmic relationship with its own frequency. This indicates a scale-free connection structure between instructions, similar to the topology of Internet. Such a pattern may offer new insights into the scaling laws of model performance and suggest ways to better understand the internal knowledge structure of datasets, potentially improving training efficiency and effectiveness. This further highlights the importance of understanding the content structure of instruction data, and underscores the value of the high-quality labeling system proposed in this work.
@@ 3 Construction of InfinityInstruct-Subject
The construction process of InfinityInstruct-Subject (InfInstruct-Sub) is illustrated in Figure 1. The process starts by collecting high-quality seed instructions. We systematically gathered all general-domain instruction datasets both manually created and automatically generated using GPT-4(Achiam et al., 2023) or ChatGPT available up to March 2024, resulting in a data pool of approximately 7 million samples. The included datasets are listed in Table 1 of the Appendix. Next, we developed an automatic labeling system based on large language models (LLMs) to analyze the distribution of the data pool. From the perspectives of coverage and depth, we selected a set of high-information seed instructions. Then, we applied an evolutionary algorithm to generate over one million new instruction samples by evolving towards greater complexity and difficulty. Building on both the seed and synthesized data, we constructed a model deficiency diagnosis system. This system identifies gaps in model capabilities and guides the targeted synthesis of new data to efficiently address those weaknesses. Finally, we implemented a strict semantic similarity–based data leakage prevention framework to detect and mitigate potential risks of data leakage throughout the construction process.
@@ 3.1 Hierarchical Multilingual Tagging System for Characterizing Instruction Content Distribution
Since LLMs may describe the same knowledge or skill using different expressions (e.g., “math calculation” vs. “mathematical calculation”), we normalize tags based on semantic similarity. Specifically, we obtain embeddings for all tags using BGE (Chen et al., 2024), and identify semantically similar tags whose cosine similarity exceeds an empirical threshold of $\lambda=0.91$. Within each group of similar tags, we retain the one with the highest frequency as the representative. To further refine the tag set, we apply DBSCAN (Ester et al., 1996) clustering with parameters $eps=0.47$ and $min\_samples=2$ to merge closely related tags. These thresholds are chosen based on empirical observations. After normalization and clustering, tags with frequency less than 100 are considered noisy and removed, following prior work (Lu et al., ). As a result, a total of 21,378 fine-grained tags are retained.
@@
Given the set of fine-grained tags, we select the top 1,000 most frequent ones, then use GPT-4 to automatically summarize them into a set of domain-level categories, so that we employ each induced category as a domain-level tag. To establish a mapping between fine-grained tags and domain-level tags, we use the prompt shown in Figure 4 along with the categorization criteria illustrated in Figure 1 of the Appendix. Based on this setup, we employ Qwen2.5-72B-Instruct (Yang et al., 2024) to generate the mappings between fine-grained and domain tags.
@@ 3.2 Informative Seed Instructions Selection
The initial data distribution largely determines the effectiveness of subsequent data synthesing, thus selecting a set of informative seed instructions is crucial for synthesizing high-quality instruction data. Given the abundance of existing instruction datasets, aimlessly synthesizing existing data is of limited value. In contrast, it is more meaningful to focus on instructions that are currently underrepresented in existing datasets (expanding coverage), or those that are relatively difficult and expose model limitations in reasoning or knowledge (enhancing depth). As shown in Figure 5, we design four criteria to select informative instructions:
1. Hard-to-Follow Instructions: We select the 50k instructions with the smallest reduction in token-level average cross-entropy loss after fine-tuning, ensuring the retained instructions are challenging to follow.
@@ 4. Undertrained Instructions:
[…]
In particular, we use Llama-2-7B-base (Touvron et al., 2023) to compute the loss values for instructions. To estimate the reduction ratio in token-level average cross-entropy loss, we fine-tune Llama-2-7B-base on a balanced instruction subset with 1,000 instructions per domain category. Table 1 presents the sources of the seed data and the number of instructions retained after filtering. By applying the above criteria sequentially, we select approximately 1.2 million informative instructions from the overall data pool.
@@ 3.3 Evolutionary Algorithm for In-Distribution Generalization
Although many instruction datasets have been synthesized, most are based on simple strategies like Self-Instruct(Wang et al., 2023), offering limited incremental value. To achieve meaningful expansion, we design a three-step evolution process that targets both coverage and depth as illustrated in Figure 6.
@@
1. Metadata-Guided Random Evolution: Based on the previously selected metadata, we adopt the Evol-Instruct method and randomly choose one of four dimensions—diversity, more reasoning steps, concretizing, or deepening, to guide each instruction’s evolution using State-of-the-Art advanced large model as synthesizer. This promotes diversity and avoids narrow expansion patterns.
2. Validation and Filtering: Evolved instructions are evaluated by an advanced large model as an AI assistant to ensure instruction-following quality. Those that fail are filtered out, ensuring the final dataset remains clean and reliable.
3. Multi-Turn Dialogue Generation: Each valid instruction is used to generate 1–4 rounds of dialogue, with the AI assistant simulating different roles.
@@ 3.4 Deficiency Diagnosis and Defect-Driven Instruction Synthesis
Despite comprehensive data collection efforts, many topics remain underrepresented, and the model may still fail to capture key knowledge and skills within instruction set, leading to capability and knowledge deficiency. While expanding the seed set or synthesizing more data via evolutionary methods can help, these approaches often produce redundant and inefficient results. To address this issue, we adopt a more targeted strategy: diagnosing model deficiencies and synthesizing instructions specifically to fill those gaps.
As illustrated in Figure 7, we first construct a diagnosis dataset $\mathcal{D}={\mathcal{D}_{i}}{i=1}^{N}$ drawn from the same distribution as the seed dataset, where each $\mathcal{D}i={x_{i},y_{i}}$ represents an input query and its corresponding reference response. We then use the fine-tuned model $M{\text{FT}}$ to generate a response $\hat{y}_{i}$ for each $x_{i}$. Next, we compare $\hat{y}_{i}$ with $y_{i}$ using an State-of-the-Art advanced large Oracle model, identifying knowledge or deficiencies in $\hat{y}_{i}$ with the prompt shown in Figure 8. Given the diagnosed issues, we then prompt the Oracle model to generate new instructions that specifically address these deficiencies using the prompt in Figure 9, to target generate instructions that remedy deficiencies within the model.
@@ 3.5 Data Leakage Prevention
Overlap between training and evaluation data can cause models to achieve inflated scores on familiar samples, leading to overestimated performance. We observe that many public instruction datasets include questions from common benchmarks like MT-Bench (Zheng et al., 2023a) and AlpacaEval (Li et al., 2023), which risks misleading evaluation. To address this issue, we use the BGE model to compute semantic similarity between evaluation queries and instructions in the training data. This process involves multi-dimensional comparisons, including semantic understanding and structural features, to ensure accurate and comprehensive matching. Based on the similarity scores, we filter out overlapping entries, effectively preventing data leakage and preserving the reliability of model evaluation.
@@ 4.1 Experimental Setup
We finetune open-source pretrained models Qwen2-7B-base and LLaMA3-8B-base on the InfInstruct-Sub dataset, and compare their performance against their respective official instruction-tuned and alignment-tuned versions. Evaluations are conducted on widely adopted LLM-judging-based benchmarks AlpacaEval 2.0 (Li et al., 2023) and Arena-Hard-V0.1 (Li et al., 2024). The reference model for AlpacaEval 2.0 is GPT-4-1106; on Arena-Hard-V0.1 is GPT-4-0314.
@@ 4.2 Experimental Results
+Self-Instruct (Wang et al., 2023) 100k 7.21 4.0
+ShareGPT (Zheng et al., 2023a) 112k 9.73 6.5
+Evol Instruct (Xu et al., 2023) 143k 8.52 5.1
+OpenHermes 1 (Teknium, 2023a) 243K 9.94 4.4
+Tulu V2 Mix (Ivison et al., 2023) 326K 9.91 5.4
+WildChat (Zhao et al., ) 652k 14.62 8.7
+OpenHermes 2.5 (Teknium, 2023b) 1M 12.89 8.2
+GenQA (University of Maryland, 2024) 6.47M 9.05 3.0
+UltraChat (Ding et al., 2023) 208k 8.29 3.6
+Magipie (Xu et al., 2024) 300k 22.66 14.9
Llama-3-8B-Instruct $>$10M 22.92 20.6
@@ 5.1 Dataset Distribution
We further compare the semantic coverage of InfInstruct-Sub with similar instruction datasets, including Alpaca(Taori et al., 2023), llm-sys(Zheng et al., 2023b), and Magpie (Xu et al., 2024). We sample 20,000 instructions from each dataset and apply the embedding and projection method same to abovementioned. As shown in Figure 12, InfInstruct-Sub exhibits a broader semantic spread, demonstrating that our strategy of selecting more low-frequency instructions as seeds effectively increases coverage in the synthesized data.
@@
To evaluate the difficulty of synthesized data, we follow the approach proposed in Magpie (Xu et al., 2024), using a large language model to assign difficulty scores to instruction samples. The difficulty is rated on a five-point scale: ’very easy’: 0, ’easy’: 1, ’medium’: 2, ’hard’: 3, ’very hard’: 4. This evaluation process is independent of the method we use for selecting high-difficulty seed instructions. Specifically, we employ Qwen-2.5-32B-Instruct as the scorer and randomly sample 1,500 instructions from each of the following datasets: Alpaca-GPT, LLM-Sys, Magpie, and InfinityInstruct-Subject.

### v2

@@ 1 Abstract
Instruction tuning has become a foundation for unlocking the capabilities of large-scale pretrained models and improving their performance on complex tasks. Thus, the construction of high-quality instruction datasets is crucial for enhancing model performance and generalizability. Although current instruction datasets have reached tens of millions of samples, models finetuned on them may still struggle with complex instruction following and tasks in rare domains. This is primarily due to limited expansion in both “coverage” (coverage of task types and knowledge areas) and “depth” (instruction complexity) of the instruction set. To address this issue, we propose a systematic instruction data construction framework, which integrates a hierarchical tagging system, an informative seed selection algorithm, an evolutionary data synthesis process, and a model deficiency diagnosis with targeted data generation. These components form an iterative closed-loop to continuously enhance the coverage and depth of instruction data. Based on this framework, we construct InfinityInstruct-Subject, a high-quality dataset containing $\sim$1.5 million instructions. Experiments on multiple foundation models and benchmark tasks demonstrate its effectiveness in improving instruction-following capabilities. Further analyses suggest that InfinityInstruct-Subject shows enlarged coverage and depth compared to comparable synthesized instruction datasets. Our work lays a theoretical and practical foundation for the efficient, continuous evolution of instruction datasets, moving from data quantity expansion to qualitative improvement.
@@ 2 Introduction
Instruction tuning serves as a cornerstone for unleashing the vast knowledge and reasoning capabilities of large pretrained models (longpre2023flan; ouyang2022training). Consequently, the construction of high-quality instruction datasets has become essential for improving model performance and generalization ability. Several efforts have constructed instruction datasets through manual annotation or automatic synthesis, with the total size of these datasets having reached tens of millions (ahmad2025opencodeinstruct; nayak2024learning). However, current large models still struggle with complex instruction-following tasks and show limitations in “rare” tasks with low frequency (qin2025incentivizing).
A main reason would be that existing works overlooked the distribution of existing data, and thus failed to efficiently enhance the “coverage and “depth of synthesized instructions, and thus limit the model performance in solving rare tasks and difficult tasks (wang2024survey; liu2024best; long2024llms). The “coverage of an instruction dataset refers to the coverage of task types and knowledge categories it covers, while the ‘depth reflects the complexity of instructions within individual tasks. Both two dimensions are essential to enlarge the capability boundaries of large models (shengyu2023instruction). Expanding the coverage improves the model’s ability to generalize across domains, while increasing depth helps enhance complex reasoning ability and compositional generalizability (sinha2024survey; zhong2025survey). However, a key challenge remains: how to design effective, controllable and interpretable instruction generation strategies that continuously expand the instruction space along both dimensions, guiding models to acquire higher-level capabilities (zhang2023survey; qian2022controllable; qin2022cold). Currently, there is a lack of unified methodology and empirical research in this area.
@@
Based on this framework, we construct the InfinityInstruct-Subject (InfInstruct-Sub) dataset, which contains $\sim$1.5 million high-quality instructions. The code and dataset are available at https://github.com/BAAI-DIPL/InfinityInstruct-Sub and https://huggingface.co/datasets/BAAI/Infinity-Instruct/tree/main/Gen. Evaluations on multiple benchmarks demonstrate its effectiveness in improving instruction-following ability. Across foundation models, our tuned models outperform their official instruction-tuned counterparts. Data analysis further reveals that, compared with comparable synthetic datasets such as Magpie (magicoder) and UltraChat (ding2023enhancing), our framework could obtain an instruction set with higher coverage and complexity. Moreover, interestingly, we observe a scaling law in the co-occurrence structure of tags of instructions: the co-occurrence frequency of a tag with others follows a negative logarithmic relationship with its own frequency. This indicates a scale-free connection structure between instructions, similar to the topology of the Internet. Such a pattern may offer new insights into the scaling laws of model performance and suggest ways to better understand the internal knowledge structure of datasets, potentially improving training efficiency and effectiveness. This further highlights the importance of understanding the content structure of instruction data, and underscores the value of the high-quality tagging system proposed in this work.
@@ 3 Construction of InfinityInstruct-Subject
The construction process of InfinityInstruct-Subject (InfInstruct-Sub) is illustrated in Figure 1. The process starts by collecting high-quality seed instructions. We systematically gathered all general-domain instruction datasets both manually created and automatically generated using GPT-4(achiam2023gpt) or ChatGPT available up to March 2024, resulting in a data pool of approximately 7 million samples. The included datasets are listed in Table 1 of the Appendix. Next, we developed an automatic tagging system based on large language models (LLMs) to analyze the distribution of the data pool. From the perspectives of coverage and depth, we selected a set of high-information seed instructions. Then, we applied an evolutionary algorithm to generate over one million new instruction samples by evolving towards greater complexity and difficulty. Building on both the seed and synthesized data, we constructed a model deficiency diagnosis system. This system identifies gaps in model capabilities and guides the targeted synthesis of new data to efficiently address those weaknesses. Finally, we implemented a strict semantic similarity–based data leakage prevention framework to detect and mitigate potential risks of data leakage throughout the construction process.
@@ 3.1 Hierarchical Multilingual Tagging System for Characterizing Instruction Content Distribution
Since LLMs may describe the same knowledge or skill using different expressions (e.g., “math calculation” vs. “mathematical calculation”), we normalize tags based on semantic similarity. Specifically, we obtain embeddings for all tags using BGE (chen2024bge), and identify semantically similar tags whose cosine similarity exceeds an empirical threshold of $\lambda=0.91$. Within each group of similar tags, we retain the one with the highest frequency as the representative. To further refine the tag set, we apply DBSCAN (ester1996density) clustering with parameters $eps=0.47$ and min_samples=2min\_samples=2 to merge closely related tags. These thresholds are chosen based on empirical observations. After normalization and clustering, tags with frequency less than 100 are considered noisy and removed, following prior work (lu2023instag). As a result, a total of 21,378 fine-grained tags are retained.
@@
Given the set of fine-grained tags, we select the top 1,000 most frequent ones, then use GPT-4 to automatically summarize them into a set of domain-level categories, so that we employ each induced category as a domain-level tag. To establish a mapping between fine-grained tags and domain-level tags, we use the prompt shown in Figure 4 along with the categorization criteria illustrated in Figure 1 of the Appendix. Based on this setup, we employ Qwen2.5-72B-Instruct (qwen2.5) to generate the mappings between fine-grained and domain tags.
@@ 3.2 Informative Seed Instructions Selection
The initial data distribution largely determines the effectiveness of subsequent data synthesis, thus selecting a set of informative seed instructions is crucial for synthesizing high-quality instruction data. Given the abundance of existing instruction datasets, aimlessly synthesizing existing data is of limited value. In contrast, it is more meaningful to focus on instructions that are currently underrepresented in existing datasets (expanding coverage), or those that are relatively difficult and expose model limitations in reasoning or knowledge (enhancing depth)(li2024quantity; shen2025long; huang2025musc; sun2024conifer). As shown in Figure 5, we design four criteria to select informative instructions:
1. Hard-to-Follow Instructions: Inspired by (li2023quantity; qin2024infobench), we select the 50k instructions with the smallest reduction in token-level average cross-entropy loss after fine-tuning, ensuring the retained instructions are challenging to follow.
@@ 4. Undertrained Instructions:
These are instructions for which the base model performs poorly with a high loss value. This may either due to either a lack of exposure or inherent difficulty. We include 200k instructions with loss greater than $\text{mean(loss)}+1.96\times\text{std(loss)}$.
In particular, we use Llama-2-7B-base (touvron2023llama) to compute the loss values for instructions. To estimate the reduction ratio in token-level average cross-entropy loss, we fine-tune Llama-2-7B-base on a balanced instruction subset with 1,000 instructions per domain category. Table 1 presents the sources of the seed data and the number of instructions retained after filtering. By applying the above criteria sequentially, we select approximately 1.2 million informative instructions from the overall data pool.
@@ 3.3 Evolutionary Algorithm for In-Distribution Generalization
Based on the previously selected informative instructions as seeds, we employ an evolutionary algorithm to synthesize new instructions, so as to ensure the performance on instruction sets similar to the seed instruction distribution, meanwhile enlarge the coverage and increasing the depth of seed instructions. Specifically, we design a three-step evolution process:
@@
1. Metadata-Guided Evolution: In this step, we adopt the widely-adopted Evol-Instruct (xu2023wizardlm) method. By randomly choose one of four dimensions—diversity, i.e., more reasoning steps, concretizing, deepening, and more diversified, to guide each instruction’s evolution using State-of-the-Art advanced large model as synthesizer.
2. Validation and Filtering: Following (tornberg2023chatgpt), evolved instructions are evaluated by an advanced large model to judge the quality, as well as the similarity between the original seed instruction.
3. Multi-Turn Dialogue Generation: Following (ding2023enhancing), each valid instruction is used to generate 1–4 rounds of dialogue, with the AI assistant simulating different roles.
@@ 3.4 Deficiency Diagnosis and Defect-Driven Instruction Synthesis
Despite comprehensive data collection efforts, many topics remain underrepresented, and the model may still fail to capture key knowledge and skills within the instruction set, leading to capability and knowledge deficiency. While expanding the seed set or synthesizing more data via evolutionary methods can help, these approaches often produce redundant and inefficient results. To address this issue, we adopt a more targeted strategy: directly diagnosing model deficiencies and synthesizing instructions specifically to fill those gaps.
As illustrated in Figure 7, we first construct a diagnosis dataset $\mathcal{D}={\mathcal{D}}_{i=1}^{N}$ drawn from the same distribution as the seed dataset, where each $\mathcal{D}_{i}=\{x_{i},y_{i}\}$ represents an input query and its corresponding reference response. We then use the fine-tuned model $M_{\text{FT}}$ to generate a response $\hat{y}_{i}$ for each $x_{i}$. Next, we compare $\hat{y}_{i}$ with $y_{i}$ using a State-of-the-Art advanced large Oracle model, identifying knowledge or deficiencies in $\hat{y}_{i}$ with the prompt shown in Figure 8. Given the diagnosed issues, we then prompt the Oracle model to generate new instructions that specifically address these deficiencies using the prompt in Figure 9, to target generate instructions that remedy deficiencies within the model.
@@ 3.5 Data Leakage Prevention
Overlap between training and evaluation data can cause models to achieve inflated scores on familiar samples, leading to overestimated performance. We observe that many public instruction datasets include questions from common benchmarks like MT-Bench (zheng2023judging) and AlpacaEval (li2023alpacaeval), which risks misleading evaluation. To address this issue, we use the BGE model to compute semantic similarity between evaluation queries and instructions in the training data. This process involves multi-dimensional comparisons, including semantic understanding and structural features, to ensure accurate and comprehensive matching. Based on the similarity scores, we filter out overlapping entries, effectively preventing data leakage and preserving the reliability of model evaluation.
@@ 4.1 Experimental Setup
We finetune open-source pretrained models Qwen2-7B-base and LLaMA3-8B-base on the InfInstruct-Sub dataset, and compare their performance against their respective official instruction-tuned and alignment-tuned versions. Evaluations are conducted on widely adopted LLM-judging-based benchmarks AlpacaEval 2.0 (li2023alpacaeval) and Arena-Hard-V0.1 (li2024live). The reference model for AlpacaEval 2.0 is GPT-4-1106; on Arena-Hard-V0.1 is GPT-4-0314.
@@ 4.2 Experimental Results
+Self-Instruct (wang2023self) 100k 7.21 4.0
+ShareGPT (zheng2023judging) 112k 9.73 6.5
+Evol Instruct (xu2023wizardlm) 143k 8.52 5.1
+OpenHermes 1 (OpenHermes1) 243K 9.94 4.4
+Tulu V2 Mix (ivison2023camels) 326K 9.91 5.4
+WildChat (zhaowildchat) 652k 14.62 8.7
+OpenHermes 2.5 (OpenHermes2) 1M 12.89 8.2
+GenQA (genqa_dataset) 6.47M 9.05 3.0
+UltraChat (ding2023enhancing) 208k 8.29 3.6
+Magipie (xu2024magpie) 300k 22.66 14.9
Llama-3-8B-Instruct >>10M 22.92 20.6
@@ 5.1 Dataset Distribution
We further compare the semantic coverage of InfInstruct-Sub with similar instruction datasets, including Alpaca(taori2023stanford), llm-sys(zheng2023lmsys), and Magpie (xu2024magpie). We sample 20,000 instructions from each dataset and apply the embedding and projection method same to abovementioned. As shown in Figure 12, InfInstruct-Sub exhibits a broader semantic spread, demonstrating that our strategy of selecting more low-frequency instructions as seeds effectively increases coverage in the synthesized data.
@@
To evaluate the difficulty of synthesized data, we follow the approach proposed in Magpie (xu2024magpie), using a large language model to assign difficulty scores to instruction samples. The difficulty is rated on a five-point scale: ’very easy’: 0, ’easy’: 1, ’medium’: 2, ’hard’: 3, ’very hard’: 4. This evaluation process is independent of the method we use for selecting high-difficulty seed instructions. Specifically, we employ Qwen-2.5-32B-Instruct as the scorer and randomly sample 1,500 instructions from each of the following datasets: Alpaca-GPT, LLM-Sys, Magpie, and InfinityInstruct-Subject.